OpenStudy (anonymous):

Need help with Calculus and Vectors Rates of Change. The time "t", in seconds, taken by an object dropped from a height of "s" metres to reach the ground is given by the formula (t= sqrt s/5). Determine the rate of change in time with respect to height when the object is 125m above the ground.

OpenStudy (anonymous):

\[t(125)=\sqrt{\frac{ 125 }{ 5 }}\] \[t(125)=5\]

OpenStudy (anonymous):

Then I have to use this formula \[t(s)= \frac{ t(s+h )- t(s)}{ h }\] lim h-->0

OpenStudy (anonymous):

lim h-->0 \[=\frac{ t(125+h)-t(125) }{ 5 }\] \[=\frac{ \sqrt{\frac{ 125+h }{ 5 }} -5}{ h }\]

OpenStudy (anonymous):

idk what to do next

OpenStudy (dumbcow):

multiply/divide by conjugate \[\frac{\sqrt{\frac{125+h}{5}} -5}{h} *\frac{\sqrt{\frac{125+h}{5}} +5}{\sqrt{\frac{125+h}{5}} +5} = \frac{\frac{125+h}{5} -25}{h(\sqrt{\frac{125+h}{5}} +5)}\]

OpenStudy (anonymous):

I got the answer thanks anyways tho

OpenStudy (dumbcow):

ok :)